Grimple, Grumple, Snarly, and Lumpy sourpusses all, wanted to divide up a giant pickle that is 30 inches long. Snarly wanted twice as much as Grimple does. Grumple wants twice as much as Snarly does but half as much as Lumpy does.
How should they divide the pickle?

Answer by mananth(12270)

let Grimple have x inches
Snarly wants twice as Grimple =2x
Grumple wants twice as much as Snarly=4x
..
lumpy gets 30-(x+2x+4x)
=30-7x
Grumple gets 1/2 (30-7x)=4x
30-7x=2*4x
30=8x+7x
30=15x
x=2 inches Grimple
Snarly gets 2x= 2*2 =4 inches
grumple= 4x= 8 inches
lumpy gets 30-7x= 30-7*2=16 inches
...
2+4+8+16 = 30 inches . the size of the giant pickle